OpenStudy (anonymous):

What is the slope of the line that passes through the points (-2, 4) and (3, 4)?

myininaya (myininaya):

m=(4-4)/(3-(-2))=0/5=0

OpenStudy (anonymous):

could you simplify?

myininaya (myininaya):

simplify 0? no